A block of mass m1 is suspended vertically with a rope connected to a block of mass m2. The block of mass m2 is on a frictionless plane inclined at theta with respect to the horizontal. If m1=2kg, the acceleration of the block is 1.96 m/s^2 downward. If m1=1kg the acceleration of the block is 1.22 m/s^2 upwards. Determine theta (angle) and m2.


1
Expert's answer
2021-11-11T12:00:37-0500

"\begin{cases}
 m_1\vec a=m_1\vec g+\vec T, \\ 
 m_2\vec a=m_2\vec g+\vec T;
\end{cases}"

"m_2(g\sin \theta±a)=m_1(g\mp a),"

"\begin{cases}
 m_2(9.81\sin\theta+1.96)=2\cdot(9.81-1.96), \\ 
 m_2(9.81\sin\theta-1.22)=1\cdot(9.81+1.22);
\end{cases}"

where

"\theta=\arcsin0.89=62.87°,"

"m_2=1.47~kg."